Lab 3: Your turn to practice Turn this in - it is a graded lab activity he Scenario: Quarterback Bill throws the ball to the right at a speed of 15 m/s. It is intercepted 45 m away by Carlos, who is running to the left at 7.5m/s. Carlos carries the ball 60 m to score. Let x0 m be the point where Bill throws the ball. Task: Represent the motion of the football through the use of multiple representations Must haves Lab questions: 1. State the total time for the entire play. 2. What is the total distance traveled by the 3. What is the total displacement of the ball? Explain how you determined this. ball? Explain how you determined this. How is this different to the total distance traveled by the ball? D inb
Answer #1

1.

Time for entire play = Time of travel for the ball from Bill to carlos + Time of travel for carlos

t = t1 + t2

d1 = distance traveled by the ball between Bill and Carlos = 45 m

v1 = velocity of the ball between Bill and Carlos =15 m/s

Time of travel for the ball from Bill to carlos is given as

t1 = d1 /v1

t1 = 45/15 = 3 sec

d2 = distance traveled by the Carlos = 60 m

v2 = velocity of Carlos =7.5 m/s

Time of travel for carlos is given as

t2 = d2 /v2

t2 = 60/7.5 = 8 sec

hence total time of travel is given as

t = t1 + t2 = 3 + 8 = 11 sec

2)

Total distance traveled by ball = distance traveled by ball from Bill to carlos + distance traveled by carlos carrying the ball =

D = d1 + d2 = 45 + 60 = 105 m

3)

Total displacement is given as

D' = d1 - d2

D' = 45 - 60

D' = - 15 m

hence the magnitude of displacement is smaller than the magnitude of the distance traveled